A piece of fabric measures 37 by 38 inches. A triangular scarf with a height of 32 inches and a base of 26 inches is cut from the fabric. How much fabric is left over?
Answer by mananth(16946)

A piece of fabric measures 37 by 38 inches. A triangular scarf with a height of 32 inches and a base of 26 inches is cut from the fabric. How much fabric is left over?
Area of fabric = 37*48 = 1776=
area of triangular piece= 1/2 32*26= 416
1776-416=1360 square inches is left over